OpenStudy (anonymous):

find the derivative of the trigonometric function… h(theta)=5 theta sec theta+theta tan theta CALCULUS HELP PLEASE!

OpenStudy (anonymous):

product rule + product rule

OpenStudy (anonymous):

do I have to find the derivative for the product rule?

OpenStudy (anonymous):

yes, product rule is a method of taking a derivative when you have a product of two functions in this case \[5\theta*\sec \theta \]is one product, while the other is\[\theta*\tan \theta \]

OpenStudy (anonymous):

Given two multiplied functions, f and g resulting in f*g the derivative is f'*g+f*g'